# ATLAS: Adaptive Trading with LLM AgentS Through Dynamic Prompt Optimization and Multi-Agent Coordination

### v1

@@ Abstract
Large language models show promise for financial decision-making, yet deploying them as autonomous trading agents raises fundamental challenges: how to adapt instructions when rewards arrive late and obscured by market noise, how to synthesize heterogeneous information streams into coherent decisions, and how to bridge the gap between model outputs and executable market actions. We present ATLAS (Adaptive Trading with LLM AgentS), a unified multi-agent framework that integrates structured information from markets, news, and corporate fundamentals to support robust trading decisions. Within ATLAS, the central trading agent operates in an order-aware action space, ensuring that outputs correspond to executable market orders rather than abstract signals. The agent can incorporate feedback while trading using Adaptive-OPRO, a novel prompt-optimization technique that dynamically adapts the prompt by incorporating real-time, stochastic feedback, leading to increasing performance over time. Across regime-specific equity studies and multiple LLM families, Adaptive-OPRO consistently outperforms fixed prompts, while reflection-based feedback fails to provide systematic gains [^1] Code will be available upon publication..
@@ 1 Introduction
Financial markets represent one of humanity’s most complex decision-making environments, requiring synthesis of vast information from technical indicators and fundamental analysis to breaking news and market sentiment. Large Language Models (LLMs) introduce new possibilities for financial decision-making through their ability to process diverse data sources and reason over complex scenarios.
From the model’s perspective, financial trading serves as an ideal testbed due to its combination of unambiguous metrics, sequential complexity, multimodal reasoning requirements, and inherent stochasticity. Unlike synthetic benchmarks, markets provide extensive historical data without simulation bias and reward genuine understanding over pattern memorization. LLMs can therefore be tasked to make decisions under uncertainty, revealing capabilities in complex reasoning He et al. (2025), market understanding Li et al. (2025), and high-risk decision-making Hung et al. (2023).
Turning these capabilities into reliable systems raises practical questions: i) how a model adapts with delayed, noisy rewards, ii) how diverse signals form coherent guidance, and iii) how decisions are expressed to align with executable market actions. Recent progress touches parts of this space, yet adaptation under delayed feedback and action space expressivity remain under-specified.
We introduce ATLAS (Adaptive Trading with LLM AgentS) to address these points. ATLAS centers on Adaptive-OPRO, a prompt optimization mechanism for sequential settings, and an order-aware decision layer that ties model outputs to execution. Structured analyses of market data, news, and fundamentals provide consistent inputs to the decision policy, while optimized instructions determine how these inputs are synthesized over time. Our experiments show that different LLMs develop distinct trading behaviors and that the combination of adaptive prompting with coordinated inputs improves decision quality over fixed prompts. Crucially, we find that reflection-based reasoning, although often assumed to be beneficial, can degrade performance in well-tuned systems, and that single-run evaluations can conceal substantial variance and lead to unstable conclusions.
These findings outline practical principles for deploying LLMs in high-stakes settings with sequential decisions under real-world uncertainty. In parallel, ATLAS improves explainability of agent decisions through explicit order specifications, structured analyst inputs, and transparent prompt evolution, supporting collaboration with financial experts and more reliable deployment.
@@ 2 Related Work
Recent work explores several LLM-based trading agents, from sentiment-driven pipelines Kirtac and Germano (2024) to coordinated, multi-component systems Zhou et al. (2025); Yang et al. (2025); Liu et al. (2023). Examples include CryptoTrade, which integrates on/off-chain signals with a reflection component Li et al. (2024), and TradingAgents, which organizes fundamental, sentiment, and technical analysts with debate mechanisms Xiao et al. (2025). Memory-centric designs such as FinMem emphasize persistent, task-specific recall Yu et al. (2023), while FINCON introduces conceptual verbal reinforcement to shape multi-agent collaboration Yu et al. (2024). Other directions incorporate learning signals Xiong et al. (2025) or mixture-of-experts routing Ding et al. (2025), and focus on document-centric analysis such as filings and earnings calls Fatouros et al. (2025). Despite advances, issues remain: manually crafted prompts with no systematic adaptation under delayed, noisy feedback, and simplified action spaces that collapse execution into directional scores. We therefore study order-level outputs (type, size, timing, price) in a simulator built for such interfaces Papadakis et al. (2025), and adopt multi-run reporting to address stochastic variability Song et al. (2025); Atil et al. (2025).
@@
Prompt optimization enhances LLM performance beyond manual tuning. Optimization by PROmpting (OPRO) treats the model as a meta-optimizer over instruction text and has shown gains on single-turn tasks with immediate feedback Yang et al. (2024). Extensions explore evolutionary search and reinforcement-style updates Guo et al. (2025); Do et al. (2024); Austin and Chartock (2024). These settings typically assume fast, unambiguous scoring and independent instances. In contrast, trading provides deferred, noisy reward signals and sequentially coupled decisions. Adaptive-OPRO adapts prompt optimization to this regime by using rolling evaluation windows and by separating static instructions from dynamic run-time content, allowing stability where consistency matters and controlled evolution where change is beneficial.
@@ 3 ATLAS Framework
ATLAS comprises three main components: (i) a Market Intelligence Pipeline, which consists of specialized agents that prepare market, news, and fundamental inputs for downstream decisions; (ii) a Decision & Execution Layer centered on a Central Trading Agent that generates and executes orders; and (iii) a feedback mechanism that collects post-execution signals and feeds them back for continuous adaptation. Within the feedback mechanism we introduce Adaptive-OPRO, an extension of the OPRO framework that dynamically edits the Central Trading Agent’s instruction prompt based on real-time, stochastic market feedback. Figure 1 provides an overview of the ATLAS framework.
3.1 Market Intelligence Pipeline
ATLAS separates information preparation from decision-making. Thus, the Market Intelligence Pipeline consists of three distinct agents, each with a specialized analyst role. Market Analyst produces multi-timescale summaries from price and volume in varying time scales (2 years, 6 months, and 3 months of history with monthly, weekly, and daily candlesticks, respectively). Within each window it computes standard indicators (e.g., moving averages, momentum, volatility bands, support/resistance) and refreshes daily, providing a consistent, noise-filtered description rather than trading signals (details in App. B). News Analyst aggregates relevant articles into structured fields (Sentiment Assessment, Key Developments, Market Relevance, Source Analysis) with optional full-text retrieval to move beyond headlines (details in App.  C.1). Fundamental Analyst extracts material changes from periodic reports and corporate events, activating infrequently to mirror reporting cycles and provide medium- to long-horizon context (details in App.  C.2).
@@ 3.1 Market Intelligence Pipeline
3.2 Decision & Execution Layer
@@
The Decision & Execution Layer determines trading actions (e.g., buying or selling a stock), executes these orders, and receives corresponding market feedback. The main decision-making component within this layer is the Central Trading Agent. Central Trading Agent consumes the structured inputs and current portfolio and emits orders that specify type (market, limit, stop), size, timing, and price levels. Orders are executed in StockSim Papadakis et al. (2025), which enforces core trading semantics and returns fills, positions, and cash for the next step. Order-level decisions clarify intent and link analytical quality to execution choices.
3.3 Adaptive-OPRO
Adaptive-OPRO updates the prompt of the Central Trading Agent using realized outcomes, targeting sequential decisions with late, noisy reward feedback. It edits only the agent’s instruction text while maintaining analyst texts, addressing three needs: credit assignment under delay (updates reflect decision sequences rather than single steps), stability under noise (avoiding short-lived fluctuations), and edit locality (preserving the run-time interface that injects analyst outputs and portfolio state).
Template separation.
The Central Trading Agent prompt is a template with two parts: static instructions (decision criteria, risk preferences, output schema) and dynamic run-time content (current Market/News/Fundamental summaries, portfolio, recent orders). Adaptive-OPRO edits only static instructions; placeholders for dynamic fields are preserved, preventing overfitting to transient data and keeping the execution path intact (Appendix G).
Windowed scoring and choice of signal.
Feedback is aggregated over windows of $K=5$ decision steps. For each window, cumulative return on investment (ROI) is computed and mapped to a score $s\in[0,100]$ following the OPRO scoring convention (Eq 1) Yang et al. (2024), so that $20\%\!\mapsto\!0$, $0\%\!\mapsto\!50$, $20\%\!\mapsto\!100$.
$s=\mathrm{clip}_{[0,100]}\big(50+250\cdot\mathrm{ROI}\big)$ (1)
The choice of $K$ and this linear scaling is carefully selected for our evaluation period and can be adjusted for other horizons or volatility regimes without change. ROI is directly aligned with the objective (capital growth) and is deterministically computed from portfolio value; when windowed and clipped it provides a low-variance, delay-aware signal that reflects the cumulative effect of a sequence of decisions rather than any single outcome.
Meta-prompted update.
At each window boundary, Adaptive-OPRO builds a prompt evolution history containing prior instruction variants and their scores. A meta-prompt presents this history to an optimizer LLM and requests four outputs: (i) a brief performance analysis that identifies likely bottlenecks, (ii) a proposed instruction update, (iii) a concise list of key improvements, and (iv) an expected impact summary. This analysis makes the rationale for edits explicit and ties changes to observed behavior, clearly showing how the policy evolves. The candidate replaces the current instruction block if the template’s placeholder set is unchanged, ensuring compatibility with the run-time injector.
@@ 3.3 Adaptive-OPRO
Claude Sonnet 4
Baseline -7.26± 2.99 -0.066± 0.030 17.59± 1.55 31.19± 7.84 13.00± 4.36
@@
Claude Sonnet 4
w/ Thinking
Baseline -4.46± 4.76 -0.043± 0.048 14.32± 4.12 11.11± 19.24 14.00± 2.65
@@
Table 1: Performance comparison between non-LLM-based and LLM-based approaches using ATLAS in volatile, declining market conditions (LLY, healthcare sector). Bold values indicate the best per model.
@@ 4 Experiments
Our study examines ATLAS along three axes: (1) Adaptation – whether sequential prompt optimization via Adaptive-OPRO improves over strong fixed prompts and over analytical reflection when feedback is delayed and noisy; (2) Component attribution – the contribution of structured inputs (Market Analyst, News Analyst, Fundamental Analyst) under different regimes; (3) Model capabilities – how backbone LLMs perform as both decision policies and prompt optimizers under Adaptive-OPRO, assessed by return and risk-adjusted performance, robustness across runs, and their ability to propose instruction updates that yield sustained improvements over windows.
4.1 Experimental Design
@@
We evaluate listed equities in three regime-specific windows that emphasize distinct conditions: bearish-volatile (LLY), sideways (XOM), and bullish (NVDA). Each window spans two months (Apr 28-Jun 28, 2025) with a daily decision interval: the agent may act once per trading day. This horizon is chosen to (i) capture multiple decision cycles without regime mixing, so adaptation reflects outcomes rather than macro shifts, and (ii) preserve complete conversation history (analyst summaries, orders, prompt-evolution logs) within the context limits of all backbones, enabling fair, auditable runs across models and ablations.
To avoid outcome-driven selection, assets were identified in advance using simple, transparent criteria (liquidity, sector diversity, representative regime behavior). ATLAS is horizon- and asset-agnostic; the same protocol applies to alternative windows or instruments, with the scoring window and scale chosen to match the volatility profile.
Models.
@@
We evaluate seven backbones spanning families, sizes, and reasoning modes: GPT-o3, GPT-o4-mini, Claude Sonnet 4 with and without thinking, LLaMA 3.3-70B, Qwen3-235B, and Qwen3-32B. Within a run, the same backbone powers all ATLAS components to avoid cross-model effects. This mix lets us examine how capability and architecture shape sequential decision making, instruction adherence, and stability, and it tests whether Adaptive-OPRO transfers across model families without per-model tuning.
Prompting strategies.
@@
We compare three strategies for the Central Trading Agent: Baseline – a strong fixed prompt produced via iterative expert prompt engineering; Reflection – a weekly reflection mechanism adapted from Li et al. (2024) that generates analytical feedback the agent must interpret; Adaptive-OPRO – our sequential prompt optimization with windowed scoring and template separation (Section 3.3). Our objective is to isolate the adaptation mechanism under identical data and execution semantics, so we use a single, transparent experimental setup rather than re-creating full external agent stacks, which differ in action-space definitions, state handling, and execution interfaces and would confound comparisons. Reflection is included as a widely used, portable form of sequential feedback, providing a focused and fair point of contrast to Adaptive-OPRO and the fixed baseline.
Non-LLM baselines.
@@
Following Li et al. (2024), we include five widely used quantitative strategies to contextualize results: Buy & Hold, MACD Wang and Kim (2018), SMA Gencay (1996), SLMA Wang and Kim (2018), and Bollinger Bands Day et al. (2023). For window-based methods, we test multiple window lengths per regime and report a strong, representative configuration for each strategy (e.g., 10-day SMA; 10/30-day SLMA). Full specifications appear in Appendix D.7.
Execution environment.
Agents interact with StockSim Papadakis et al. (2025) through an order-level interface. The simulator enforces core trading semantics (submitting orders with price levels and sizes, updating positions and cash, and reporting fills) and produces a complete audit trail of orders, executions, and portfolio state. This setup moves beyond generic directional views and evaluates whether a policy can translate analysis into concrete, executable actions. We intentionally abstract fine-grained market microstructure to keep runs comparable and to isolate the decision policy. The objective of our evaluation is order-aware reasoning and coherent execution under standard semantics, not microstructure modeling. Fixing these mechanics makes results deterministic and auditable across models and ablations.
@@
Sharpe Ratio (SR): Risk-adjusted return metric calculated as: $\frac{\mu-r_{f}}{\sigma}$, where $\mu$ is mean daily return, $\sigma$ is daily return standard deviation, and $r_{f}$ is the risk-free rate (set to 0 following Li et al. (2024)). Higher values indicate better returns per unit of risk.
Maximum Drawdown (DD): The worst peak-to-trough decline in portfolio value: $\max_{t\in[0,T]}\left(\max_{s\in[0,t]}V_{s}-V_{t}\right)/\max_{s\in[0,t]}V_{s}$, where $V_{t}$ is portfolio value at time $t$. This measures the largest loss from any historical high point, capturing the strategy’s downside risk and stress tolerance.
Win Rate: Percentage of closed (i.e., completed) trades that are profitable:
@@
Protocol and statistics.
To address the stochasticity of LLM outputs, each configuration is run three times and we report results as mean $\pm$ standard deviation. This multiple-run protocol captures variability and helps distinguish genuine capability differences from random fluctuations. Together with the reported distributional and execution diagnostics, it surfaces behavioral aspects that single-run summaries may miss.
Model Prompting XOM NVDA
@@
Claude Sonnet 4
Baseline -4.49± 4.22 -0.134± 0.114 7.71± 1.06 13.43± 8.62 0.180± 0.121 5.52± 3.96
@@
Claude Sonnet 4
w/ Thinking
Baseline -0.99± 0.80 -0.039± 0.020 7.75± 1.00 12.52± 2.47 0.175± 0.030 5.03± 1.53
@@
Table 2: Combined performance table across two markets: XOM (range-bound) and NVDA (bullish). Includes ROI, SR, and $D$. Bold values indicate the best results per model. Full results are available in Appendix E.

### v2

@@ Abstract
Large language models (LLMs) offer promising capabilities for financial decision-making, yet their deployment in sequential trading settings faces two key challenges: synthesizing heterogeneous information sources and adapting agent behavior under delayed and noisy reward signals. We address these challenges by introducing ATLAS (Adaptive Trading with LLM AgentS), a unified agentic framework for systematic integration of market data, financial news, and corporate fundamentals, and Adaptive-OPRO, a novel prompt optimization method that dynamically updates agent instructions using real-time stochastic feedback. We evaluate our approach across regime-specific equity trading scenarios and multiple LLM families. Results demonstrate that Adaptive-OPRO consistently outperforms existing methods, particularly in highly volatile regimes. Moreover, our analysis reveals that increased information availability does not necessarily translate to improved performance, highlighting the importance of careful modality integration in noisy market environments. [^1] Code will be released upon publication.
@@ 1 Introduction
Financial markets represent one of humanity’s most complex decision-making environments, requiring synthesis of vast information from technical indicators and fundamental analysis to breaking news and market sentiment. LLMs introduce new possibilities for financial decision-making through their ability to process diverse data sources and reason over complex scenarios.
From the model’s perspective, financial trading serves as an ideal testbed: it combines unambiguous metrics, sequential complexity, multimodal reasoning requirements, and inherent stochasticity. Unlike synthetic benchmarks, markets provide extensive historical data without simulation bias and reward genuine understanding over pattern memorization. LLMs can therefore be tasked to make decisions under uncertainty, revealing capabilities in complex reasoning He et al. (2025), market understanding Li et al. (2025), and high-risk decision-making Hung et al. (2023).
Despite this potential, stock market decision-making introduces inherent challenges beyond stochasticity. Decisions require synthesizing heterogeneous signals such as price dynamics, market conditions, and firm-specific developments into coherent actions. Moreover, in high-stakes financial environments where capital is continuously at risk, static decision policies are insufficient; decision patterns must be revised by incorporating market feedback as it unfolds, enabling continuous behavioral adaptation.
Consequently, turning LLM capabilities into reliable trading systems raises two key queries: (i) how diverse signals are synthesized into coherent guidance, and (ii) how models adapt their behavior through continuous market interaction. While recent work explores these issues partly, their systematic study in realistic trading settings is limited.
In this work, (i) we propose ATLAS, a multi-agent framework that provides a foundational structure for experimentation in LLM-based stock market decision-making; (ii) we introduce Adaptive-OPRO, a prompt optimization mechanism for sequential settings that supports behavioral adaptation through ongoing market interaction and achieves state-of-the-art performance across multiple market regimes and LLM families. Through extensive regime-aware evaluations, we show that additional input modalities are not uniformly beneficial and depend critically on market conditions.
@@ 2 Related Work
Recent work explores several LLM-based trading agents, from sentiment-driven pipelines Kirtac and Germano (2024) to coordinated, multi-component systems Zhou et al. (2025); Yang et al. (2025); Liu et al. (2023). Examples include CryptoTrade, which integrates on/off-chain signals with reflection Li et al. (2024), and TradingAgents, which coordinates specialized roles via structured debate and synthesis Xiao et al. (2025). Memory-centric designs such as FinMem emphasize persistent, task-specific recall Yu et al. (2023), while FINCON introduces conceptual verbal reinforcement to shape multi-agent collaboration Yu et al. (2024). Other works incorporate learning signals Xiong et al. (2025) or mixture-of-experts routing Ding et al. (2025), and focus on document-centric analysis such as filings and earnings calls Fatouros et al. (2025). However, key limitations persist: prompts are usually hand-crafted even when feedback is delayed and noisy, and many setups collapse execution into directional scores. Our approach pairs a prompt-tuning component with order-level evaluation (type, size, timing, price) in a simulator built for such interfaces Papadakis et al. (2025), using multi-run reporting to account for stochastic variability Song et al. (2025); Atil et al. (2025).
@@
Prompt optimization enhances LLM performance beyond manual tuning. Optimization by PROmpting (OPRO) treats the model as a meta-optimizer over instruction text and has shown gains on single-turn tasks with immediate feedback Yang et al. (2024). Extensions explore evolutionary search and reinforcement-style updates Guo et al. (2025); Do et al. (2024); Austin and Chartock (2024). These settings typically assume fast, unambiguous scoring and independent instances. In contrast, trading provides deferred, noisy reward signals and sequentially coupled decisions. Our Adaptive-OPRO adapts prompt optimization to this regime by using rolling evaluation windows and by separating static instructions from dynamic run-time content, allowing stability where consistency matters and controlled evolution where change is beneficial.
@@ 3 ATLAS Framework
ATLAS comprises three main components: (i) a Market Intelligence Pipeline, which consists of specialized agents that prepare market, news, and fundamental inputs for downstream decisions; (ii) a Decision & Execution Layer centered on a Central Trading Agent that generates and executes orders; and (iii) a feedback mechanism that collects post-execution signals and feeds them back for continuous adaptation. Within the feedback mechanism we incorporate Adaptive-OPRO, an extension of the OPRO framework that dynamically edits the Central Trading Agent’s instruction prompt based on real-time, stochastic market feedback. Figure 1 provides an overview of the ATLAS framework.
Market Intelligence Pipeline.
ATLAS separates information preparation from decision-making. The Market Intelligence Pipeline consists of three specialized agents, each with a distinct analyst role. Market Analyst produces multi-timescale summaries from price and volume in varying time scales (2 years, 6 months, and 3 months of history with monthly, weekly, and daily candlesticks, respectively). Within each window it computes standard indicators (e.g., moving averages, momentum, volatility bands, support/resistance) and refreshes daily, providing a consistent, noise-filtered description rather than trading signals (details in App. B). News Analyst aggregates relevant articles into structured fields (Sentiment Assessment, Key Developments, Market Relevance, Source Analysis) with optional full-text retrieval to move beyond headlines (details in App.  C.1). Fundamental Analyst extracts material changes from periodic reports and corporate events, activating infrequently to mirror reporting cycles and provide medium- to long-horizon context (details in App.  C.2).
@@
Decision & Execution Layer.
@@
This layer determines trading actions (e.g. buying or selling a stock), executes these orders, and receives corresponding market feedback. The main decision-making component within this layer is the Central Trading Agent (CTA). This agent consumes the structured inputs and current portfolio and emits orders that specify type (market, limit, stop), size, timing, and price levels. Orders are executed in StockSim Papadakis et al. (2025), which enforces core trading semantics and returns fills, positions, and cash for the next step. Order-level decisions clarify intent and link analytical quality to execution choices.
Feedback Mechanism.
This mechanism defines how information derived from market outcomes is incorporated into the agent’s future decisions. It may be entirely absent, resulting in a static agent that follows a fixed policy, or it may be enabled to support adaptation based on observed performance. In general, the mechanism processes signals such as returns or behavioral outcomes from past decisions and uses them to influence subsequent actions. Implementations can range from simple feedback summaries to more structured optimization approaches, such as reflection-based methods Li et al. (2024). In the following section, we describe Adaptive-OPRO, a prompt optimization technique that leverages market feedback to iteratively refine the agent’s decision-making process.
4 Adaptive-OPRO
Adaptive-OPRO is a sequential prompt-optimization procedure that improves an agent’s instruction prompt using delayed, noisy performance feedback. It generalizes OPRO to interactive settings where decisions are temporally coupled and rewards arrive after multiple steps. The core idea is to treat instruction text as the optimized object and to update it periodically via a learned update mechanism (implemented by an optimizer LLM), while keeping the agent’s run-time inputs and interfaces stable.
Optimized object, state, and round inputs.
Adaptive-OPRO maintains a current instruction prompt $P_{t}$ for a target agent that acts over time, along with an optimization history $\mathcal{H}=\{(P_{i},s_{i})\}_{i<t}$ storing past prompt variants and their scores. At the end of each evaluation window, it constructs an optimizer query with the following inputs: (i) a meta-prompt $M$ that specifies the optimizer’s role and constraints, (ii) $\mathcal{H}$ (or a compact summary thereof), and (iii) a summary of the agent’s recent interaction outcomes together with a scalar performance score $s_{t}$ for $P_{t}$. An update rule $U$, implemented using an optimizer LLM, produces a revised prompt $P_{t+1}=U(M,\mathcal{H},s_{t},\text{summary})$. The output of each round is an updated instruction prompt that governs subsequent agent decisions.
Stability via template separation.
In sequential systems, prompt updates can inadvertently break the run-time interface (e.g., input placeholders, output schemas) or overfit to transient observations. Adaptive-OPRO therefore separates the target agent prompt into: (a) static instructions (policy, priorities, constraints, formatting requirements), and (b) dynamic run-time content injected at execution time (state, observations, tool outputs, recent actions). Only the static instruction block is editable; all placeholders and the run-time injection format are held fixed. This enforces edit locality: updates can change how the agent reasons and decides, but cannot change what information it receives nor the interface it must comply with.
Windowed evaluation under delayed feedback.
To address credit assignment and reduce variance, Adaptive-OPRO evaluates prompts over rolling windows of $K$ decision steps. After each window, the system computes a scalar performance score $s$ from outcomes observed during that window (e.g. task success, reward, utility, risk-adjusted return). The choice of $K$ and scoring function is task-dependent; the only requirement is that $s$ provides consistent ordering to compare prompt variants.
Meta-prompted update rule.
At each window, Adaptive-OPRO forms the optimizer query from $M$, $\mathcal{H}$, and the recent outcome summary/score, and applies $U$ to generate a candidate $P_{t+1}$. The optimizer is instructed to: (i) diagnose likely failure modes of the current prompt, (ii) propose a revised instruction prompt $P_{t+1}$, (iii) summarize the concrete changes made, and (iv) state the expected behavioral impact. The candidate is accepted only if it preserves the template (e.g., required placeholders and output schema). The accepted prompt is appended to history with its subsequent window score, enabling iterative improvement.
ATLAS instantiation.
In ATLAS, Adaptive-OPRO is applied to the Central Trading Agent’s instruction prompt (i.e., the static instruction block of the decision policy). Dynamic run-time content corresponds to the daily injected analyst summaries, portfolio state, and recent executions, which are kept fixed by construction (Appendix I). For scoring, we aggregate portfolio performance over K=5K{=}5 trading days to reduce noise and capture delayed effects of sequential decisions, then map cumulative ROI to a bounded OPRO-style score $s\in[0,100]$ via linear scaling and clipping:
$s=\mathrm{clip}_{[0,100]}\big(50+250\cdot\mathrm{ROI}\big),$ (1)
so that $20\%\!\mapsto\!0$, $0\%\!\mapsto\!50$, $20\%\!\mapsto\!100$. This yields a stable, delay-aware signal while limiting the impact of outlier windows; the optimizer is restricted to instruction edits that preserve ATLAS’s execution interface.
@@ 4 Adaptive-OPRO
Claude Sonnet 4 Baseline -7.26± 2.99 -0.066± 0.030 17.59± 1.55 31.19± 7.84 13.00± 4.36
@@
Claude Sonnet 4 w/ Thinking Baseline -4.46± 4.76 -0.043± 0.048 14.32± 4.12 11.11± 19.24 14.00± 2.65
@@
Table 1: Performance comparison between non-LLM-based and LLM-based approaches using ATLAS in volatile, declining market conditions. Bold values indicate the best per model.
@@ 5 Experiments
Our study examines ATLAS along three axes: (1) Adaptation – whether sequential prompt optimization via Adaptive-OPRO improves over well-tuned static prompts and over analytical reflection when feedback is delayed and noisy; (2) Component attribution – contribution of structured inputs (Market Analyst, News Analyst, Fundamental Analyst) under different regimes; (3) Model capabilities – performance of backbone LLMs as both decision policies and prompt optimizers under Adaptive-OPRO, assessed by return and risk-adjusted performance, robustness across runs, and their ability to propose instruction updates for sustained improvements over windows.
5.1 Experimental Setup
@@
Specifically, we evaluate stock market decision-making across three distinct market regimes: a bearish-volatile regime characterized by declining prices and elevated uncertainty, a sideways regime marked by range-bound price dynamics and limited directional trends, and a bullish regime defined by sustained upward momentum and comparatively favorable risk-return conditions. Each window spans two months (Apr 28-Jun 28, 2025) with a daily decision interval: the agent may act once per trading day. This horizon is chosen to (i) capture multiple decision cycles without regime mixing, so adaptation reflects outcomes rather than macro shifts, and (ii) preserve complete conversation history (analyst summaries, orders, prompt-evolution logs) within the context limits of all backbones, enabling fair, auditable runs across models and ablations. More details in App. D.1
The experimental setup, including the evaluation method, metrics, regime partitioning, and evaluation horizon, follows Li et al. (2024), ensuring methodological consistency and fair comparison across settings. We explicitly account for LLM stochasticity by running each configuration three times and reporting mean $\pm$ standard deviation, distinguishing systematic performance differences from randomness rather than single-run variability.
@@
Models. We evaluate seven backbones spanning families, sizes, and reasoning modes: GPT-o3, GPT-o4-mini, Claude Sonnet 4 with and without thinking, LLaMA 3.3-70B, Qwen3-235B, and Qwen3-32B. Each run uses a single backbone for all ATLAS components and Adaptive-OPRO, isolating how model capacity and architecture affect sequential behavior, instruction adherence, stability, and cross-family transfer without per-model tuning.
@@
Prompting strategies. We compare three strategies for the Central Trading Agent: Baseline – a fixed instruction prompt obtained via iterative expert prompt engineering; Reflection Li et al. (2024) – a weekly reflection mechanism that summarizes recent trajectories into high-level feedback that the agent must interpret; Adaptive-OPRO – our sequential prompt optimization with windowed scoring and template separation (Section 4). Our goal is to isolate the adaptation mechanism under identical data and execution semantics. We therefore evaluate all methods within a single, transparent setup rather than re-implementing full external agent stacks, which differ in action spaces, state representations, and execution interfaces. We include reflection as a widely used and portable form of sequential feedback, providing a focused comparison to Adaptive-OPRO and the fixed baseline.
@@
Non-LLM baselines. Following Li et al. (2024), we include five widely used quantitative strategies to contextualize results: Buy & Hold, MACD Wang and Kim (2018), SMA Gencay (1996), SLMA Wang and Kim (2018), and Bollinger Bands Day et al. (2023). For window-based methods, we test multiple window lengths per regime and report a strong, representative configuration for each strategy (e.g., 10-day SMA; 10/30-day SLMA). Full specifications in App.  D.8.
Execution environment. Agents interact with StockSim Papadakis et al. (2025) via an order-level action space, requiring CTAs to submit fully executable orders (type, side, size, price). Compared to signal- or position-level formulations common in prior LLM trading studies Li et al. (2024); Xiao et al. (2025), this enforces execution feasibility (cash, inventory, validity) while yielding a complete audit trail of orders, fills, and portfolio states. Consistent with standard offline evaluation, we abstract away market microstructure and assume deterministic execution, ensuring observed differences stem from decision policies rather than execution frictions.
@@
Sharpe Ratio (SR): Risk-adjusted return metric calculated as: $\frac{\mu-r_{f}}{\sigma}$, where $\mu$ is mean daily return, $\sigma$ is daily return standard deviation, and $r_{f}$ is the risk-free rate (set to 0 as in Li et al. (2024)).
Maximum Drawdown (DD): The worst peak-to-trough decline in portfolio value: $\max_{t\in[0,T]}\left(\max_{s\in[0,t]}V_{s}-V_{t}\right)/\max_{s\in[0,t]}V_{s}$, where $V_{t}$ is portfolio value at time $t$. This measures the largest loss from any historical high, reflecting downside risk and stress tolerance.
@@
Model Prompting Sideways Market Bullish Market
@@
Claude Sonnet 4 Baseline -4.49± 4.22 -0.134± 0.114 7.71± 1.06 13.43± 8.62 0.180± 0.121 5.52± 3.96
@@
Claude Sonnet 4 w/ Thinking Baseline -0.99± 0.80 -0.039± 0.020 7.75± 1.00 12.52± 2.47 0.175± 0.030 5.03± 1.53
@@
Table 2: Combined performance table across two markets: range-bound (sideways) and bullish market. Includes ROI, SR, and $D$. Bold values indicate the best results per model. Full results are available in Appendix E.